[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Mar 27, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Mar 27 13:31:50 UTC 2017


HSU Shift Log for Mon Mar 27 2017 09:31:44 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: Ridges is still down. The backup work station hangs up  
periodically if you try to do any complicated work. Sawfish is not working  
properly. The only machine that appears to be working properly is Tilefish  
(Atlantic side with the big monitor).

AWIPS-2/NCP issues: 6z GFS data very late, over one hour. This has been a  
problem for almost a week now. Something about the dataflow recently has  
been quite slow and has delaying that model from coming in on time.

Other Techincal issues: ATT Trouble Ticket 732402127 on our circuit is  
still open, with very loud hum still coming from the hotline. I tried to  
contact them and they were going to do testing later today.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170327/7dc0dd5f/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list